A large group of students took a test in Physics and the final grades have a mean of 80 and a standard deviation of 7. 97.98% of the students received a score on the test that was greater than what grade?
Answer by Theo(13342)

the mean was 80.
the standard deviation was 7.
assuming a normal distribution, the critical z-score would be -2.049635638.
to find the raw score, you would use the following formula.
z = (x-m)/s
z = -2.049635638..., which you would round off to -2.05.
x = what you want to find.
m = 80.
s = 7
formula becomes -2.05 = (x-80)/7
solve for x to get x = -2.05 * 7 + 80.
this results in a raw score of 65.65.
assuming the rounded numbers are accurate enough, then you can assume that 97.98% of the scores on the test were greater than or equal to 65.65.

pictorially, this looks like .....

the shaded area in the graph is the part that applies.

$$$